A spacecraft of surface area of 60m² with a temperature of 600K if the radiation that it receives from the sun is equivalent to a temperature in space of 120K, find the net rate of loss of heat energy from the spacecraft


1
Expert's answer
2021-08-12T08:56:28-0400

The power loss per unit area (radiant emittance) is given by the Stefan–Boltzmann law:


"j_{loss} = \sigma T_1^4"

where "\sigma = 5.67\times10^{-8}\dfrac{W}{m^2\cdot K^4}" is the Stefan–Boltzmann constant, and "T_1 = 600K" is the temperature of the spacecraft.

The power gain per unit area is given similarly:


"j_{gain} = \sigma T_2^4"

where "T_2 = 120K" is the temperature of the space. Thus, net rate loss from all area ("A = 60m^2" ) is:


"P = A(j_{loss} - j_{gain}) = A\sigma (T_1^4 - T_2^4)"

"P = 60\cdot 5.67\times10^{-8}\cdot (600^4 - 120^4) = 4.4\times 10^{5}W"

Answer. "4.4\times 10^5W".